-(x - 6) + 3 = -12
What do I do with the negative sign in front of the parentheses ?

Found 2 solutions by lenny460, MathTherapy:
Answer by lenny460(1073)

-(x - 6) + 3 = -12
What do I do with the negative sign in front of the parentheses ?

x = 21

Answer by MathTherapy(10806)


-(x - 6) + 3 = -12
What do I do with the negative sign in front of the parentheses ?

An "invisible" 1 exists after the - (negative), or between the negative and the open parenthesis.
Therefore, - (x - 6) + 3 = - 12 is the same as: - 1(x - 6) + 3 = - 12
You now need to distribute the - 1, and solve for x